OpenStudy (anonymous):

solve -4J-(-5) __________=13 -7

OpenStudy (anonymous):

-4J-(-5) __________=13 multiply by -7 on both sides -7 -4J-(-5) =-91 simplify left hand side -4J + 5 =-91 subtract 5 from both sides -4J =-96 divide by -4 on both sides J =24

jhonyy9 (jhonyy9):

-4+5=-91 1=-91

jhonyy9 (jhonyy9):

how you have thought it solving for x or y or z or for what unknown ?

jhonyy9 (jhonyy9):

oh sorry this is 4j all right -4j+5=-91 -4j=-96 j=24